OpenStudy (anonymous):

Suppose a bag contains 7 blue chips and 3 gold chips. What is the probability of randomly choosing a blue chip, not replacing it, and then randomly choosing a gold chip? A. 7/30 B. 1/15 C. 1/30 D. 7/15

OpenStudy (anonymous):

7/30

OpenStudy (anonymous):

i was thinking C.

OpenStudy (anonymous):

The chance of drawing a blue chip on the first draw is 7/10 because there are 7 blue chips out of 10 in the bag. Now after that blue chip is removed, the chance of getting a gold chip is 3/9 (or 1/3) because there are 3 gold chips out of 9 in the bag. The probability of getting both those events is the product: P(blue then gold) = P(first is blue) x P(second is gold) = 7/10 x 3/9 = 7/10 x 1/3 = 7/30 ≈ 0.2333 (23.33%)

OpenStudy (anonymous):

thats how you know its simple actually

OpenStudy (anonymous):

wow ok

OpenStudy (anonymous):

Okay thanks, i got the fraction part right but i wasnt sure on where to go from there. I appreciate the help.